A pendulum of length 98.1000cm. The mass at the end of the pendulum is 0.2kg. Calculate the period and frequency of pendulum.
1
Expert's answer
2018-08-16T10:05:09-0400
The period of pendulum oscillations
T=2π√(l/g)
=2×3.14×√(0.981/9.81)=2.0 s
The frequency of pendulum oscillations
f=1/T
=1/2.0=0.5 Hz